Home‎ > ‎Reports‎ > ‎Crew104‎ > ‎

day04


Crew 104 Log Book for April 26, 2011

Photos of the day





Commander's Check-In Report

Jon Clarke Reporting

Crew Physical Status: Good

Time Departed/Returned from EVA: Drill partly left at 9:20 am, returned at 1 pm after completion of first core.  Commenced drilling second core and 2:00 om, completed second core and packing up MARTE drill for the night at 6:30 pm

First science DVA departed at 9:20 to carry out baseline biological and biochemical sampling sampling at drill site and an undisturbed site some 200 m away, returned at 10:30.  Second science EVA at 12:00 noon to collect first core, returned by 12:30 pm.  Third science EVA departed at 2:00 to sample sites from last year, returned at 4 pm

Brief Narrative of Field Mission Results:  Drilling successful, meeting minimum crew objectives. 24 core drilled, 19 in first run, 5 in second.  Successfully collected samples for contamination at drill site and control site.  Successfully collected soil sample from upper Dakota Formation for repeat analysis.

EVA Data/Interpretations: Pending

Engineering/Hab Maintenance: No outstanding issues

Report Transmission Schedule:  Journalist report sent, engineering and commanders report pending 

Plans for Tomorrow:  Continue drilling, re-sample two further sites, sample analysis.

Inventory: No issues

Support Requested: None at this stage

Miscellaneous: None


Commander's Report

Jon Clarke Reporting

This has been a most successful day. After a wild night, weather wise, a sunny, cold, windy day dawned.  The MARTE drill was fired up and succeeded in drilling 24 cm of core. Conditions were quite arduous for the crew, with 30 kph wind and significant wind chill.  However, no rain is forecast for the next five days to we expect to be able to drill to 2 m.  Even the core recovery we achieved today met the minimum success criteria for the crew.

Also impart for the biological analysis of the core is to know the degree of surface contamination, so the biologists were sampling the drill site and a control site several hundred meters away to quantify this issue.  The sites were geologically documented as well.

Carol and Jon also  revisited a site from Crew 83 in 2009 that showed anomalous amino acid levels.  This site was in the upper Dakota Formation and repeat samples were taken to test these results. 

We are all getting to the routine and timing our activities better.  the crew is sleeping better and enjoying  good food.  Chicken Biryana tonight!

Tomorrow we intend to keep drilling and carry out further repeat and infill soil samples.



Engineering Report

David Willson Reporting

Generator/Electricity:

Kitty (Used? Yes/No): Used Continuously and not stopped

Honey (Diesel Generator) (Used? Yes/No): Not Used

Wendy (Gas Generator) (Used? Yes/No): Not Used

Battery Charge Level: Not Checked

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

Generator operating very well continuously and not stopped


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 35% in Tank

Propane Tank (Full/0.75/0.5/0.25/Empty): 35%

Gasoline Tank (Full/0.75/0.5/0.25/Empty): N/A

Notes/Comments:


Oil:

Kitty Oil Quantity (# of Quarts in storage): Not Checked as used continuously

Honey Oil Quantity (# of Quarts in storage): Not Checked

ATV Oil Quantity (# of Quarts in storage): Not Checked

Car Oil Quantity (# of Quarts in storage): Not Checked


Water Status:

Outside Potable Water Tank Level (inches from bottom): Full

Trailer Potable Water Tank Level (inches from bottom): 15 cm

Hab Potable Water Tank Level (inches from bottom): 50%

Potable Water Meter Reading: 455380

In to GreenHab Meter Reading: N/A

Notes/Comments: Green Hab not operational but water is being recycled


GreenHab:

Condition of plants in Tank 1: 5%

Condition of plants in Tank 2: Full

Condition of plants in Tank 3: 15% Full

Greenhab Notes/Comments: Green house not used

Observatory:

Telescope Used? (Yes/No): No

Observatory Notes/Comments:

Transportation:

Hab Car (NOT RENTAL): Operating

Used (yes/no): NO

Oil Checked (yes/no):

Tire Status:

Odometer at end of day:

Notes/Comments on Hab Car:


ATV 1 (Opportunity): Operating

Used: y/n NO

Oil Checked: y/n

Fuel Consumed:

Tire Status:

Comments:


ATV 2 (Spirit):

Used: y/n NO

Oil Checked: y/n

Fuel Consumed:

Tire Status:

Comments:


ATV 3 (Viking I):

Used: y/n NO

Oil Checked: y/n

Fuel Consumed:

Tire Status:

Comments:


ATV 4 (Viking II):

Used: y/n NO

Oil Checked: y/n

Fuel Consumed:

Tire Status:

Comments:


Heating and Ventilation:

Thermostat setting upstairs (Farenheit): A/C not operating 61 degrees F

Actual temperature upstairs (Farenheit):

Thermostat setting downstairs (Farenheit): 55 degrees F

Actual temperature downstairs (Farenheit): 50 degrees F

Computers/Networking Infrastructure: Operating

Notes/comments on internet/computers?: All OK


General Engineering Notes/Comments:


Chef's Report


Sarah Thompson Reporting

Day 4: Chicken Biryani


Ingredients

5 cups Basmati rice with lentils
10 cups water
1 tin coconut milk
2 tins chicken in water
3 chicken stock cubes
1 small jar of ghee
1 jar of Patak's Madras curry paste
2 cups dried peas
1 cup dried brocolli
1 tin mixed vegetables

Method

1. Chop the chicken so that all the pieces are approx 0.5" across
2. Add the chicken, water, rice, half the jar of curry paste, half the jar of ghee to a large stock pot. Mix well. Cover tightly and bring to the boil, turning down head as necessary, for about 20 minutes.
3. Add the other half of the ghee, the other half of the curry paste, the drained tin of mixed vegetables, the dried brocolli and the dried peas and the tin of coconut milk, with enough water to reconstitute the dried ingredients. Cover and bring to the boil, turning down to a simmer for 45 minutes to 1 hour.

Total prep time: about 1 hour

Comments:

This is a pretty traditional Biryani made from ingredients I managed to scrounge up in the Hab. Again, I got good comments, though I should have cooked the dried peas longer (reflected in the above -- I doubled the cooking time in the recipe).

Journalist Report

Sarah Thompson Reporting

More of the same

At the time of writing, Tuesday is working out much the same as yesterday in terms of outline, though there is a funny happening I'd like to share, with permission. The MDRS hab is designed for a crew of 6, though there are actually 7 of us. Consequentially, we either need to double up in one of the rooms, or someone needs to sleep in a tent. Carol was particularly keen on camping out, so she took one for the team, so to speak. She picked a particularly idyllic spot that just happened to be right in the middle of a dry stream bed.

We had a lot of rain last night. Um, a lot of rain.

The photo is the aftermath of Carol cleaning out something like 5 gallons of mud from her tent. Looks like the doubling up option is going to work out more reliably from now on.